# Risk Twin: Real-time Risk Visualization and Control for Structural Systems

### v1

@@ Abstract
Digital twinning in structural engineering is a rapidly evolving technology that aims to eliminate the gap between physical systems and their digital models through real-time sensing, visualization, and control techniques. Although digital twins can offer dynamic insights into physical systems, their accuracy is inevitably compromised by uncertainties in sensing, modeling, simulation, and controlling. This paper proposes a specialized digital twin formulation, named Risk Twin, designed for real-time risk visualization and risk-informed control of structural systems. Integrating structural reliability and Bayesian inference methods with digital twining techniques, Risk Twin can analyze and visualize the reliability indices for structural components in real-time. To facilitate real-time inference and reliability updating, a “simulation-free” scheme is proposed, leveraging precomputed quantities prepared during an offline phase for rapid inference in the online phase. Proof-of-concept numerical and real-world Risk Twins are constructed to showcase the proposed concepts.
@@ 1 Introduction
Building on existing progress of integrating probabilistic analysis within the digital twinning technology, this study aims to further advance their application in risk assessment and control of civil structures. Specifically, this paper proposes (i) the concept of Risk Shadow as a digital reflection of structural risk; (ii) an efficient “simulation-free” method to update the Risk Shadow; and (iii) real-world benchmark Risk Twin systems. In the benchmark experiments, the bidirectional interaction between the physical and digital systems is demonstrated.
The paper is structured as follows: Section 2 introduces the basic concepts of Risk Twin. Section 4 presents implementation details for two benchmark Risk Twin systems. Additionally, demonstration videos for these Risk Twin systems are provided for further reference. Consequently, conclusive remarks are drawn in Section 5.
@@ 2 Risk Twin
If a digital twin operates with deterministic parameters, it will classify the state of a system as either “safe” or “failure.” However, uncertainties are prevalent in digital twin systems due to variabilities in sensing and modeling, hindering an exact reflection of the true state of a system. Risk twin is a probabilistic digital twin designed to provide real-time quantification, visualization, and control of risks for the modeled system. This section begins with a general introduction to the framework of Risk Twin and then delves into the detailed developments of its core components.
@@ 2.1 General framework
Figure 1: Framework of the Risk Twin. The workflow begins with sensors collecting data from the physical twin. This data is then used to update the probabilistic model of basic random variables. The updated probabilistic model is propagated into a reliability analysis module to obtain reliability indices for system components. Subsequently, the Risk Shadow, displaying these reliability indices, is presented to users. Users can then decide whether to execute control operations upon perceiving the Risk Shadow. Once a control command is received, actuators, governed by an algorithm with operational cost constraints, act on the physical system, completing the full loop of information flows between the physical and digital systems. To address the computational challenge of real-time inference and reliability updating, the operation of risk twin is decomposed into offline and online phases. During the offline phase, computationally intensive models are simulated to prepare datasets and quantities for rapid, simulation-free inference in the online phase.
@@ 5 Conclusions
This paper introduces and demonstrates the concept of Risk Twin, a digital twin system designed for real-time reliability visualization and reliability-informed control. The computational challenges of statistical inference and reliability updating are addressed by decomposing the service life of risk twins into offline and online phases. During the offline phase, computationally intensive models are simulated to prepare datasets and quantities for rapid, simulation-free inference in the online phase. To illustrate the proposed concept, two benchmark risk twin experiments are conducted using commonly available tools and materials. The first experiment demonstrates real-time Bayesian inference for quantifying the uncertainty in the position and magnitude of an external force on a simply supported plate. The second experiment showcases the risk shadow and human-risk shadow interactions using a cantilever beam controlled by a mechanical arm. Future research could explore the application and adaptation of risk twin technology to real-world civil structures and infrastructures, such as the real-time reliability-based control of wind turbines.
@@ Acknowledgement
This work is partly supported by the New Faculties’ Basic Research Capability Enhancement Program through No.2-9-2022-010 at China University of Geosciences (Beijing). Dr. Zeyu Wang appreciates the instructive comments from Prof. Daniel Straub at Technical University of Munich(TUM), the assistance of Yi Ren at Qualcomm(Shanghai) for his support in sensor connecting and Zhuoqun Gao for her meticulous care. These supports are really appreciated. Any opinions, findings, and conclusions expressed in this paper are those of the authors, and do not necessarily reflect the views of the sponsors.
Appendix A Implementation details of the mechanical arm control

### v2

@@ Abstract
Digital twinning in structural engineering is a rapidly evolving technology that aims to eliminate the gap between physical systems and their digital models through real-time sensing, visualization, and control techniques. Although Digital Twins can offer dynamic insights into physical systems, their accuracy is inevitably compromised by uncertainties in sensing, modeling, simulation, and control. This paper proposes a specialized Digital Twin formulation, named Risk Twin, designed for real-time risk visualization and risk-informed control of structural systems. Integrating structural reliability and Bayesian inference methods with Digital Twinning techniques, Risk Twin can analyze and visualize the reliability indices for structural components in real-time. To facilitate real-time inference and reliability updating, a simulation-free scheme is proposed. This scheme leverages precomputed quantities prepared during an offline phase for rapid inference in the online phase. Proof-of-concept numerical and real-world Risk Twins are constructed to showcase the proposed concepts.
@@ 1 Introduction
Building on existing progress of integrating probabilistic analysis within the Digital Twinning technology, this study aims to further advance their application in risk assessment and control of civil structures. Specifically, this paper proposes (i) the concept of Risk Shadow as a digital reflection of structural risk; (ii) an efficient simulation-free method to update the Risk Shadow; and (iii) real-world and numerical benchmark Risk Twin systems. In the benchmark experiments, the bidirectional interaction between the physical and digital systems is demonstrated.
The paper is structured as follows: Section 2 introduces the basic concepts of Risk Twin. Section 3 presents implementation details for two benchmark Risk Twin systems. Moreover, simulation of wind turbine RT is elaborated in Section 4. Additionally, demonstration videos for these Risk Twin systems are provided for further reference. Consequently, conclusive remarks are drawn in Section 5.
@@ 2 Risk Twin
If a Digital Twin operates with deterministic parameters, it will classify the state of a system as either safe or failure. However, uncertainties are prevalent in Digital Twin systems due to variabilities in sensing and modeling, hindering an exact reflection of the true state of a system. Risk Twin is a probabilistic Digital Twin designed to provide real-time quantification, visualization, and control of risks for the modeled system. This section begins with a general introduction to the framework of Risk Twin and then delves into the detailed developments of its core components.
@@ 2.1 General framework
Figure 1: Workflow of the Risk Twin. The workflow begins with sensors collecting data from the physical twin. This data is then used to update the probabilistic model of basic random variables. The updated probabilistic model is propagated into a reliability analysis module to obtain reliability indices for system components. Subsequently, the Risk Shadow, displaying these reliability indices, is presented to users. Users can then decide whether to execute control operations upon perceiving the Risk Shadow. Once a control command is received, actuators, governed by an algorithm with operational cost constraints, act on the physical system, completing the full loop of information flows between the physical and digital systems. To address the computational challenge of real-time inference and reliability updating, the operation of Risk Twin is decomposed into offline and online phases. During the offline phase, computationally intensive models are simulated to prepare datasets and quantities for rapid, simulation-free inference in the online phase.
@@ 3.2.3 Human-Risk Shadow interaction
4 Risk Twin of a Wind Turbine
4.1 Computational model of a wind turbine
We demonstrate the application of Risk Twin using a simulation model of a wind turbine, where parameters and computational methodologies are partially referred to [36]. The proposed RT framework updates and visualizes the reliability indexes of the structural components under three limit states in real-time. Moreover, the uncertainty parameters primarily originate from wind speed, wind direction, sensor equipment, and the load-bearing capacity of the components. As shown in Figure 9, a horizontal-axis wind turbine (HAWT) consists of four key components: a tubular tower, blades, rotor hub, and nacelle. The total height of the tubular tower, $H$, is $80$ meters, including the height of the hub and nacelle. The base diameter of the tower, $D$, is $10$ meters, tapering linearly to an upper diameter $d$ of $4$ meters, with a wall thickness of $0.02$ meters. As documented in [36], the variation of the local chord of the blade along the rotor radius is presented in Table 4. The wind turbine is equipped with three blades, each with a mass of $4000$ kg, assuming the center of mass is at the midpoint of the blade. The angular acceleration due to the friction between hub and blade, is inversely proportional to the angular velocity with the coefficient $k_{f}=0.01$. To prevent the blades from rotating excessively fast and inducing damage, a brake system is installed at the rotor hub. This system activates when the blade’s rotational speed reaches a maximum value of $\omega_{max}=1.8$ rad/sec. Once the brake system is activated, the angular acceleration of the blade drops to zero, preventing further increases in angular velocity. We treat the simulated wind turbine as a physical twin, serving as the basis to demonstrate the application of Risk Twin in wind farms. A demonstration video of this simulation model can be accessed through the link: https://youtu.be/2iTJ_FDZXrg.
Figure 9: A wind turbine structure subject to wind loads: (a) Conceptual illustration of the four core components in a wind turbine and (b) the angular definitions of yaw angle, $\theta_{Y}$, wind speed, $V_{W}$, and wind direction, $\theta_{W}$. This paper considers only the deviation between $\theta_{Y}$ and $\theta_{W}$, without addressing the effects of upwind or downwind.
Table 4: Geometry of the wind turbine blade: local chord length $c$ and the corresponding radius $r$.
$r$ [m] 4.5 6.5 8.5 10.5 12.5 14.5 16.5 18.5 20.3
$c$ [m] 1.63 1.540 1.420 1.294 1.163 1.026 0.881 0.705 0.265
Wind turbines operate by harnessing wind force to drive the rotation of the blades, thereby generating mechanical energy in the rotor. This mechanical energy is then converted into electrical energy by the generator located in the nacelle. Therefore, the aerodynamic properties that facilitate the generation of mechanical energy are crucial. Figure 10(a) illustrates the mechanical analysis model of a blade with a pitch angle under operational conditions, subjected to wind speed $V_{0}$ perpendicular to the rotor plane, and the rationale for computing the relative velocity $V_{rel}$, where $V_{0}=V_{W}\sin{\left|{\theta_{w}-\theta_{Y}}\right|}$. $V_{rel}$ is decomposed into the perpendicular components $(1-a)V_{0}$ and $(1+a^{\prime})\omega r$:
$V_{rel}\sin{\theta_{rel}}=V_{0}(1-a)\,,$ (27)
and
$V_{rel}\cos{\theta_{rel}}=V_{0}(1+a^{\prime})\,,$ (28)
where $\theta_{rel}$ denotes the angle between the rotor plane and $V_{rel}$, $a$ denotes the axial induction factor due to the existing vortex system, $a^{\prime}$ is the tangential induction factor, $r$ denotes the radial length of the target position on the blade, and $\omega$ is the rotational speed of the rotor. It follows that $V_{rel}$ can be computed as:
$V_{rel}=\sqrt{(V_{0}(1-a))^{2}+(\omega r(1+a^{\prime}))^{2}}\,,$ (29)
and
$\tan{\theta_{rel}}=\frac{(1-a)V_{0}}{(1+a^{\prime})\omega r}\,.$ (30)
Figure 10: Aerodynamics of wind turbine blades: (a) Relative velocity analysis considering vortex effects and (b) Structural analysis of lift and drag forces per unit length.
The effect of the vortex system can be negligible, indicating that $a=a^{\prime}=0$. Moreover, let $\theta_{att}=\theta_{rel}-\theta_{p}$ be the local angle of attack of the wind. According to the aerodynamics of wind turbines, the lift $L$ and drag force $D$ per unit length caused by the wind can be respectively calculated by:
$L=\frac{1}{2}\rho V_{rel}^{2}cC_{l}\,,$ (31)
$D=\frac{1}{2}\rho V_{rel}^{2}cC_{d}\,,$ (32)
where $\rho$ denotes the air density, with $\rho=1.29\,\text{kg/m}^{3}$, and $c$ denotes the local chord of the blade. The coefficients $C_{l}$ and $C_{d}$ are the lift and drag coefficients, which are parameterized by $\theta_{att}$. The relationship between $C_{l}$, $C_{d}$, and $\theta_{att}$ is illustrated in Figure 11(a). It should be noted that $C_{d}$ can be significantly large if $\theta_{att}<0$. However, the more critical forces are those perpendicular to the rotor plane, denoted by $P_{N}$, and those along the rotor plane direction, denoted by $P_{T}$. $P_{N}$ acts along the blade, making the rotor highly susceptible to flapwise moments, potentially causing damage at the blade-hub connection. Additionally, $P_{T}$ drives the rotor blade’s rotation, serving as the primary source of power for the wind turbine. However, excessive $P_{T}$ can generate an overly large shaft moment, which can lead to destructive bending moments in the hub and nacelle structure. Figure 10(b) illustrates the geometric relationships among $V_{rel}$, $L$, $D$, $P_{N}$, and $P_{T}$, expressed as:
$P_{N}=L\cos{\theta_{rel}}+D\sin{\theta_{rel}}\,,$ (33)
$P_{T}=L\sin{\theta_{rel}}-D\cos{\theta_{rel}}\,.$ (34)
Consequently, the power of the wind turbine can be computed by:
$P=\frac{1}{2}\rho{V_{0}}^{3}AC_{p}(\lambda)\,,$ (35)
where $A=\pi R^{2}$ is the rotor area, $\lambda=\omega R/V_{0}$ denotes the tip-speed ratio, and $C_{p}(\lambda)$ represents the power coefficient parameterized by $\lambda$. Figure 11(b) illustrates the relationship between $C_{p}$ and $\lambda$. The coefficients $a$ and $a^{\prime}$ can be estimated through the classical Blade Element Momentum (BEM) method, expressed as:
$a=\frac{1}{\frac{4{\sin}^{2}{\theta_{rel}}}{\sigma C_{n}}+1}\,,$ (36)
and
$a^{\prime}=\frac{1}{\frac{4\sin{\theta_{rel}}\cos{\theta_{rel}}}{\sigma C_{t}}-1}\,,$ (37)
where $\sigma$ denotes the fraction of the annular area in the control volume in BEM, expressed as:
$\sigma=\frac{cB}{2\pi r}$ (38)
where $B=3$ is the number of blades. $C_{n}$ and $C_{t}$ in Eqs. (36) and (37) can be estimated by:
$C_{n}=C_{l}\cos{\theta_{rel}}+C_{d}\sin{\theta_{rel}}\,,$ (39)
and
$C_{t}=C_{l}\sin{\theta_{rel}}-C_{d}\cos{\theta_{rel}}\,.$ (40)
Figure 11: Relationships between model parameters: (a) $C_{l}$ and $C_{d}$ versus $\theta_{att}$ and (b) $C_{p}$ versus $\lambda$.
Generally, the final values of $a$ and $a^{\prime}$ are obtained through an iterative process. First, one initializes $a$ and $a^{\prime}$ and compute $\theta_{rel}$ and $\theta_{att}$ to obtain the initial values of $C_{l}$ and $C_{d}$. Next, $C_{n}$ and $C_{t}$ are computed to update $a$ and $a^{\prime}$ until the results converge. The final values of $a$ and $a^{\prime}$ can be corrected according to correction factors such as Prandtl’s tip loss factor and Glauert correction. As shown in Figure 12, the BEM discretizes the blade into $N$ elements, where $F_{T}$ denotes the tangential forces for each element. Moreover, it should be noted that the trend of force in normal direction $F_{N}$ also aligns with $F_{T}$. $F_{T}$ and $F_{N}$ can be computed based on $P_{T}$ and $P_{N}$ through the integration of all the elements along blade. Consequently, the shaft torque, $M_{shaft}$, can be calculated via the BEM method:
$M_{shaft}=\sum_{i=1}^{N-1}\left(\frac{p_{T,i+1}-p_{T,i}}{3(r_{i+1}-r_{i})}(r_{i+1}^{3}-r_{i}^{3})+\frac{p_{T,i}r_{i+1}-p_{T,i+1}r_{i}}{2(r_{i+1}-r_{i})}(r_{i+1}^{2}-r_{i}^{2})\right)\,,\$ (41)
where $r_{i}$ denotes the radial length of the $i$-th element, and $p_{T,i}$ is the force per unit length tangential to the rotor plane of the $i$-th element. The rotational speed of the rotor can be computed by:
$\dot{\omega}=\frac{M_{shaft}}{BI_{b}}-k_{f}\omega$ (42)
where $\omega$ and $\dot{\omega}$ represent the rotor’s angular speed and acceleration, respectively, and $I_{b}$ denotes the blade’s moment of inertia about the hub.
Figure 12: Illustration of discretization of elements in blade through BEM method.
4.2 Structural failures of the wind turbine
The wind turbine has a probability of encountering extreme wind loads during operation, which poses a risk of structural failure. We consider the failure modes of the blade, rotor hub, and tower. As shown in Figure 13, blade failure primarily arises from the flapwise direction moment, hub failure is mainly due to excessive shaft moment from the brake system, and tower failure occurs when the maximum stress exceeds the threshold value. Therefore, we define three limit state functions.
Figure 13: Conceptual illustration of the three failure modes of the wind turbine: (a) blade failure due to flapwise moment, (b) hub failure due to shaft moment, and (c) tower failure due to stress.
The limit-state function for blade failure is defined as the flapwise bending moment exceeding a threshold, expressed as [37]:
$g_{blade}=M_{flap}^{thr}-\left|{M_{flap}}\right|\,,$ (43)
where
$M_{flap}=\sum_{i=1}^{N-1}\left(\frac{p_{N,i+1}-p_{N,i}}{3(r_{i+1}-r_{i})}(r_{i+1}^{3}-r_{i}^{3})+\frac{p_{N,i}r_{i+1}-p_{N,i+1}r_{i}}{2(r_{i+1}-r_{i})}(r_{i+1}^{2}-r_{i}^{2})\right)\,.$ (44)
The limit-state function for the rotor hub failure is defined as the shaft moment exceeding a threshold is defined as follows:
$g_{hub}=M_{shaft}^{thr}-\left|{M_{shaft}}\right|\,,$ (45)
where
$M_{shaft}=M_{shaft}-BIk_{f}{\omega}\,.$ (46)
The tower is modeled as a bending beam subjected to a point load $F_{N}$ at the top and a distributed wind pressure that increases with height along the tower. The limit state function is defined as the maximum stress induced by these forces exceeding a threshold, expressed as:
$g_{tower}=\sigma_{tower}^{thr}-\max_{h}\sigma_{tower}\,,$ (47)
and
$\sigma_{tower}(h)=\sqrt{\sigma_{1}^{2}(h)+\sigma_{2}^{2}(h)}\,,$ (48)
where $\sigma_{tower}(h)$ denotes the stress of the tower at the height $h$, $\sigma_{1}$ denotes the stress caused by normal force $F_{N}$ and $\sigma_{2}$ is the stress induced by wind pressure. $\sigma_{1}$ and $\sigma_{2}$ can be computed as:
$\sigma_{i}(h)=\frac{M_{i}(h)D_{h}(h)}{2I_{t}},\ i=1,2\,,$ (49)
where $D_{h}$ denotes the diameter of the tower at $h$, $I_{t}$ is the moment of inertia of the cross-section, $M_{1}$ denotes the equivalent moment caused by the normal force $F_{N}$ expressed by:
$M_{1}(h)=F_{N}(H-h)\,,$ (50)
and $M_{2}$ can be computed by:
[…]
where $\eta$ is the gust factor that considers the ratio between maxima and mean of the wind velocity, and the wind velocity at the height $h$, $V_{h}(z)$, can be computed as:
$V_{h}(z)=V_{w}(z/H)^{0.15}\,.$ (52)
4.3 Risk Twin of the wind turbine
We consider the uncertainties arising from the thresholds of the limit state functions, wind speed and direction, and their measurements, as summarized in Table 5. The wind speed $V_{w}$ and wind direction $\theta_{w}$ are modeled by homogeneous Gaussian processes. Specifically, the wind speed is modeled by the following equation:
$V_{w}(t)=\Psi(t,\kappa_{w}),$ (53)
where $t$ denotes the operation time of wind turbine and $\Psi$ denotes a Gaussian process. The wind direction, $\theta_{w}$, is modeled by another Gaussian process expressed by:
$\theta_{w}(t)=\Theta(t,\kappa_{w}),$ (54)
where $\Theta$ is a Gaussian process. The kernel/correlation functions of both Gaussian processes are modeled by:
$\kappa_{w}(t_{1},t_{2})=\sigma_{f_{s,d}}^{2}\exp{\left(\frac{-(t_{1}-t_{2})^{2}}{2\tau}\right)}\,,$ (55)
where $\sigma_{f_{s,d}}^{2}$ denotes the process variance of wind speed or direction with $\sigma_{f_{s}}^{2}=1\ m/s$ and $\sigma_{f_{d}}^{2}=1^{\circ}$, and $\tau$ is the correlation length scale of the process. In this paper, we set $\sigma_{f}=1$ and $\tau=1$. Figure 14 shows a random realization of the wind speed and direction processes, where the black solid line is the random realization and red dashed line represents mean value. We assume sensors to measure the wind speed and direction are installed at the nacelle. The measurement errors $\phi_{v}$ and $\phi_{\theta}$ for wind speed and direction, respectively, follow Gaussian distributions with parameters listed in Table 5.
Figure 14: A random realization of the wind speed and wind direction.
Table 5: Random variables of wind turbine.
Random Variable Distribution Type Mean Standard Deviation
$M_{flap}^{thr}$ Lognormal $2\times 10^{5}(N\cdot m)$ $1\times 10^{4}(N\cdot m)$
$M_{shaft}^{thr}$ Lognormal $2\times 10^{5}(N\cdot m)$ $1\times 10^{4}(N\cdot m)$
$\sigma_{tower}^{thr}$ Lognormal $7.5\times 10^{6}(N/m^{2})$ $7.5\times 10^{5}(N/m^{2})$
$V_{w}$ Gaussian Process Time-Variant 1.5m/s
$\theta_{w}$ Gaussian Process Time-Variant $3^{\circ}$
$\phi_{v}$ Normal 0 0.5 m/s
$\phi_{\theta}$ Normal 0 $3^{\circ}$
We conduct two simulation experiments: (i) the forward experiment, which shows the information flow from the simulated physical model to the digital model, where the data collected from sensors are used to build the Risk Shadow, and (ii) the inverse experiment, which shows the information flow from the digital model to the physical model, where optimization of the yaw angle $\theta_{Y}$ and pitch angle $\theta_{p}$ is performed. For both experiments, the total simulation time is $400$ seconds, with initial parameters $\theta_{Y}=0^{\circ}$ and $\theta_{p}=5^{\circ}$. In the inverse experiment, $\theta_{Y}$ and $\theta_{p}\in[0,5^{\circ}]$ are adaptively tuned, with specific steps detailed in Algorithm 2 in A.
Figure 15: Risk Twin of simulated wind turbine: (a) simulated physical twin and (b) Risk Twin.
[…]
Figure 16 shows snapshots at $269$ seconds from the inverse experiment. Figure 16(a) depicts the Risk Twin based on fixed parameters, while Figure 16(b) represents the one with optimized parameters. It is seen that the optimization effectively mitigates the potential risks to the wind turbine under intense wind load conditions. Specifically, $\theta_{Y}$ was optimized from $0^{\circ}$ to $70.5^{\circ}$, and $\theta_{p}$ was optimized from $5^{\circ}$ to $29.5^{\circ}$. Additionally, the reliability index $\beta_{blade}$ was increased from $2.9$ to $13.4$, and $\beta_{tower}$ was increased from $2.6$ to $3.2$. The figure also shows a trade-off between power generation and structural reliability.
Figure 16: Risk Twin of simulated wind turbine: (a) fixed parameters and (b) optimized parameters.
Figure 17 compares the reliability indices and energy generation of the wind turbine with and without dynamic yaw and pitch angle controls, i.e., the inverse flow from the digital model to the physical model. The black solid line represents the simulation with fixed yaw and pitch angles, while the blue dashed line represents the results with dynamic angle controls. Several observations can be made: (i) The proposed Risk Twin can enhance the power generation efficiency of the wind turbine during low wind speeds. (ii) The Risk Twin can help the structure avoid potential failures during high wind speeds. (iii) It seeks a balance between structural safety and power generation.
Figure 17: Performance of wind turbine with optimization: (a) posterior reliability index of the blade, (b) posterior reliability index of the hub, (c) posterior reliability index of the tower, (d) real-time power, and (e) produced energy. The red solid line denotes the reliability threshold. During the period from 250 seconds to 350 seconds, when wind speeds are high, the control system of the Risk Twin seeks to reduce the structural risk of the blade, hub, and tower. During the period from 50 seconds to 250 seconds, when wind speeds are low, the control system seeks to increase wind power generation.
@@ 5 Conclusions
This paper introduces and demonstrates the concept of Risk Twin, a Digital Twin system designed for real-time reliability visualization and reliability-informed control. The computational challenges of statistical inference and reliability updating are addressed by decomposing the service life of Risk Twins into offline and online phases. During the offline phase, computationally intensive models are simulated to prepare datasets and quantities for rapid, simulation-free inference in the online phase. To illustrate the proposed concept, two benchmark physical experiments and one simulation experiment are conducted. The first physical experiment demonstrates real-time Bayesian inference for quantifying the uncertainty in the position and magnitude of an external force on a simply supported plate. The second physical experiment showcases the Risk Shadow and Human-Risk Shadow interactions using a cantilever beam controlled by a mechanical arm. The final numerical experiment showcases the bi-directional information flow between physical and digital models using a simulated wind turbine. Future research could explore the application and adaptation of Risk Twin technology to real-world civil structures and infrastructures.
@@ Acknowledgement
The first author is partly supported by the New Faculties’ Basic Research Capability Enhancement Program through No. 2-9-2022-010 at China University of Geosciences (Beijing). The first author appreciates the assistance of Yi Ren at Qualcomm (Shanghai) for his support in sensor connectivity and Zhuoqun Gao for her meticulous care. The authors appreciate the constructive comments from Prof. Daniel Straub at the Technical University of Munich. Any opinions, findings, and conclusions expressed in this paper are those of the authors and do not necessarily reflect the views of the sponsors.
Appendix A Implementation details of the mechanical arm and wind turbine control